by the Systel lu lllU Uu also negative. FOR PRACTICE 6.1 external A cylinder and piston assembly (defined as the system) is warmed by an flame. The contents of the cylinder expand, doing work on the surroundings by push ing the piston outward against the external pressure. If the system absorbs 559 J of heat and does 488 J of work during the expansion, what is the change in internal energy (AE) for the system?
Answer #1

Apply First law of thermo:

Q - W = dE (for closed system) i.e. the cylinder + piston

Q = 559 J ( positive since it is absorbed)

W = 488 J done by the system to surrounding, so we assing this to be postive)

now...

from

Q - W = dE

dE = Q - W = 559 -488 J = 71 J

the energy change for the system is then only 71 J